OpenStudy (brittyarellano):

What is the way you could find the "exact" value of the following sum of logarithms: log 3000 + log(1/3) a.) It doesn't have an exact value, but instead a value approximated with a decimal number b.) Use the Quotient Property of Logarithms c.) Find its equivalent value, log 1000 d.) Convert it to its equivalent exponential form

OpenStudy (brittyarellano):

i think it is A but i am not 100%sure

jimthompson5910 (jim_thompson5910):

hint: look at this page. Specifically at rule 1 http://www.purplemath.com/modules/logrules.htm

OpenStudy (mrs.ambrose614):

Did I help your girlfriend?????

OpenStudy (brittyarellano):

the rule that multiplication inside the log can be turned into addition outside the log?

jimthompson5910 (jim_thompson5910):

Yes, so using that rule \[\Large \log\left({\color{red}{A}}\right)+\log\left({\color{blue}{B}}\right) = \log\left({\color{red}{A}}*{\color{blue}{B}}\right)\] \[\Large \log\left({\color{red}{3000}}\right)+\log\left({\color{blue}{\frac{1}{3}}}\right) = \log\left({\color{red}{3000}}*{\color{blue}{\frac{1}{3}}}\right)\] \[\Large \log\left({\color{red}{3000}}\right)+\log\left({\color{blue}{\frac{1}{3}}}\right) = \log\left(1000\right)\] Do you see how to finish up?

OpenStudy (brittyarellano):

no I'm not sure I get it

jimthompson5910 (jim_thompson5910):

where are you stuck? which step?

OpenStudy (mathmale):

Hint: log a = log b = log ab (which jim has already pointed out). log 5 + log (1/2) = log (5/2). Please give this homework problem a try.